OpenStudy (theviper):

A solid is in the shape of a cone standing on a hemisphere with both their radii being equal to \(1cm\) & the height of the cone is equal to its radius. Find the volume of the solid in terms of \(\color{red}{\pi}\).

OpenStudy (theviper):

@experimentX :)

OpenStudy (experimentx):

do you know how to find the volume of cone??

OpenStudy (theviper):

yes \(\large{\dfrac{1}{3}\pi r^2h}\)

OpenStudy (experimentx):



OpenStudy (theviper):

\(\large{\dfrac{2}{3}\pi r^3=\dfrac{2\pi}{3}(1)^3=\dfrac{2\pi}{3}}\)

OpenStudy (experimentx):

okay ... put the value (for cone) r=h=1 ... and add those two.

OpenStudy (theviper):

\(\large{\dfrac{1}{3}\pi r^2h=\dfrac{1}{3}\pi (1)^2\times1=\dfrac{\pi}{3}}\)

OpenStudy (theviper):

now \(\large{\dfrac{2\pi}{3}+\dfrac{\pi}{3}=\dfrac{3\pi}{3}=\pi^3}\) ??

OpenStudy (theviper):

oh yes thanx :)